OpenStudy (anonymous):

Decide whether the pair of lines is parallel, perpendicular, or neither. 9x + 3y = 12 and 12x + 4y = 17

OpenStudy (anonymous):

well you can put the two in slope intercept form and decide from the slopes

OpenStudy (anonymous):

parallel the coefficient is -3 in front of x

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

set them up in y=mx+b form

OpenStudy (anonymous):

parallel!

OpenStudy (anonymous):

3y=-9x+12 y=-3x+4, m=-3 4y=-12x+17 y=-3x+17/4, m=-3 Parallel, like johnnybreuer said.

OpenStudy (anonymous):

sorry its parallel \[\frac{a1}{a2}=\frac{b1}{b2} \neq \frac{c1}{c2}\]